### v2

# Unveiling Entity-Level Unlearning for Large Language Models: A Comprehensive Analysis

Abstract
Large language model unlearning has garnered increasing attention due to its potential to address security and privacy concerns, leading to extensive research in the field. However, much of this research has concentrated on instance-level unlearning, specifically targeting the removal of predefined instances containing sensitive content. This focus has left a significant gap in the exploration of full entity-level unlearning, which is critical in real-world scenarios such as copyright protection. To this end, we propose a novel task of Entity-level unlearning, which aims to erase entity-related knowledge from the target model completely. To thoroughly investigate this task, we systematically evaluate trending unlearning algorithms, revealing that current methods struggle to achieve effective entity-level unlearning. Then, we further explore the factors that influence the performance of the unlearning algorithms, identifying that knowledge coverage and the size of the forget set play pivotal roles. Notably, our analysis also uncovers that entities introduced through fine-tuning are more vulnerable to unlearning than pre-trained entities. These findings collectively offer valuable insights for advancing entity-level unlearning for LLMs.
Unveiling Entity-Level Unlearning for Large Language Models:
A Comprehensive Analysis
Weitao Ma  Xiaocheng Feng♣♠thanks:     means Corresponding Author  Weihong Zhong  Lei Huang Yangfan Ye  Xiachong Feng  Bing Qin♣♠ Harbin Institute of Technology Peng Cheng Laboratory The University of Hong Kong {wtma,xcfeng,whzhong,lhuang,yfye,qinb}@ir.hit.edu.cn, fengxc@hku.hk
1 Introduction
Large Language Models (LLMs) (Achiam et al., 2023; Touvron et al., 2023a, b; Meta, 2024) pre-trained on extensive corpora have achieved significant success in knowledge-intensive tasks (Kamalloo et al., 2023; Seegmiller et al., 2024). However, undesirable data exists with training data, such as toxic texts (Lu et al., 2022), privacy content (Liu et al., 2024a) and copyrighted information (Karamolegkou et al., 2023). Such data has raised security and legal concerns, hindering the practical application of LLMs (Yao et al., 2024; Das et al., 2024). To tackle this, Machine Unlearning (Zhang et al., 2023; Lu et al., 2024; Bhardwaj et al., 2024) has gradually been applied to LLMs due to its effectiveness and cost-efficiency. These refined techniques, now known as LLM Unlearning (Yao et al., 2023; Liu et al., 2024b, c), have become a mainstream approach for removing undesirable knowledge from the model by applying post-hoc modifications to target models.
Figure 1: The comparison of the entity-level unlearning and instance unlearning processes. The knowledge covered by the purple background represents the target set, and the knowledge covered by the green background represents the forget set.
Towards this direction, existing work has conducted extensive research. However, most of these efforts focus on Instance-level Unlearning tasks, which address isolated sensitive content (Li et al., 2024; Zhang et al., 2024; Ji et al., 2024), while neglecting the deletion of entire entities, which is crucial in many real-world scenarios, such as removing ‘Harry Potter’ for copyright protection Eldan and Russinovich (2024). To address this gap, we formally define a novel task of Entity-level Unlearning. As illustrated in Figure 1, existing instance-level unlearning tasks focus on removing predefined facts by applying unlearning algorithms to a forget set that contains the specific information to be erased. In contrast, the entity-level unlearning task seeks to eliminate an entire entity, including all the entity-related knowledge within the model. However, due to the inaccessibility of the original training corpora, it is challenging to obtain such a target set, which contains all knowledge associated with the target entity. Consequently, entity-level unlearning requires an additional step to construct a forget set before applying unlearning algorithms, where there is frequently a discrepancy between the constructed forget set and the target set. To this end, we divide the entity-level unlearning task into a two-stage framework, comprising forget set construction and unlearning execution. This process seeks to delete the target set by erasing the construted forget set, which involves more complex procedures and presents greater challenges than instance-level unlearning.
In this work, we conduct a comprehensive analysis of entity-level unlearning for LLMs, structured around a sequence of focused steps: Firstly, we begin by formally defining and setting the task, which involves injecting pseudo-entity knowledge into the model to establish both the target model and the target set. Secondly, we systematically evaluate the current unlearning algorithms on this task. Specifically, within the framework, we introduce an accessible knowledge-probing method to identify entity-related knowledge as the forget sets, and then compare the performance of the unlearning algorithm across different forget sets. Our experiments reveal that existing unlearning algorithms can not generalize well to the entity-level unlearning task. Additionally, we observe that the design of forget set significantly influences the effectiveness of the unlearning algorithm. Thirdly, inspired by the above, we further investigate how various constructed forget sets impact algorithm performance. Our results indicate that unlearning with a forget set that has higher knowledge coverage relative to the target set leads to more effective unlearning. Although merely increasing the size of the forget set may enhance knowledge coverage, it would compromise the model’s generalization ability, making it an ineffective way to improve performance. Moreover, our experiments surprisingly reveal that entities introduced through fine-tuning are more vulnerable to unlearning compared to pre-trained entities, highlighting the need for more robust methods for injecting target knowledge to simulate pre-trained knowledge better.
Through extensive experiments, our study provides valuable insights for future research. We encourage future work to explore more precise knowledge-probing techniques and develop targeted unlearning algorithms to improve the preformance of entity deletion.
2 Entity-level Unlearning
2.1 Task Definition and Setting
The objective of the entity-level unlearning task is to remove an entire entity from the target model completely. However, the inaccessible training data makes it challenging to obtain comprehensive information about the entity within the model. Therefore, we define the task as removing the entire entity from the target model by deleting only a subset of entity-associated knowledge.
The entity-level unlearning task can be formalized as follows: Given a target entity O, the target model parametrized by $\theta_{t}$ is required to forget a target set $S_{T}$, which contains all knowledge related to the entity O in the model, by applying unlearning methods $\textsc{H}(\cdot)$ on a forget set $S_{F}$, which contains part of the knowledge of the entity. All of the aforementioned sets consist of pieces of knowledge, $S=\{(x_{i},y_{i})\}_{i=1}^{N}$. The unlearning process can be expressed as follows:
$\theta_{t+1}\leftarrow\textsc{H}(\theta_{t},S_{F})$ (1)
To precisely assess the deletion effect of the target entity, the evaluation for entity-level unlearning task $\textsc{E}(\cdot)$ should be conducted on the target set $S_{T}$:
$Score_{forget}=\textsc{E}(\theta_{t+1},S_{T})$ (2)
A significant challenge in evaluating the effectiveness of entity-level unlearning posed by the inability to access all their training data of LLMs is obtaining the target set $S_{T}$ used to evaluate the unlearned models. To address this, we simulate entity-level unlearning scenarios following the TOFU (Maini et al., 2024), which fine-tunes the LLMs using a fictitious author dataset. The dataset ensures that the LLM has no prior exposure to these authors during previous training phases. Thus, the fine-tuning dataset encompasses all the knowledge about the entities, making it suitable to serve as the target set $S_{T}$.
2.2 Entity-level Unlearning Framework
In light of the definition of the entity-level unlearning task, we propose a two-stage framework for the task, consisting of: 1) Forget Set Construction and 2) Unlearning Execution. Building on this framework, we design entity-level unlearning methods based on trending unlearning algorithms.
2.2.1 Forget Set Construction
The first step of the entity-level unlearning task is constructing a forget set, which is both critical and challenging. In practical scenarios, the forget set often would be generated solely based on the target model and the names of the entities involved. Following this setting, we propose a simple and effective constructing baseline to probe the entity-related question-answer (QA) pairs within the target model as a forget set. More details can be found in Appendix A.
Specifically, we first prompt the target models to self-generate entity-related questions according to their internal knowledge, inspired by Weller et al. (2023). Note that only those non-repetitive questions containing the entities’ names will be kept to further ensure the focus remains on the corresponding entities. Next, we acquire the answers from the target model using greedy decoding, which ensures that the answers possess a relatively high generation probability. Finally, we conduct a self-verification process, where the model repeatedly evaluates each QA pair, retaining only the pairs for which the model consistently agrees with its responses. These selected pairs serve as valid candidates for forget sets of unlearning techniques.
This approach offers a straightforward knowledge extraction method for acquiring the forget set. We then assess the influence of the forget set’s quality on unlearning methods through a manual replacement analysis.
2.2.2 Unlearning Execution
After constructing the forget set, the next step is to apply the unlearning algorithms to it. In the absence of algorithms specifically designed for entity-level unlearning, we select five representative algorithms for evaluation. Among them is Gradient Ascent (Grad. Ascent), which reduces the likelihood of the answers to achieve unlearning. Other methods introduce additional constraints on a retain set, which contains knowledge that should be preserved to minimize the damage to the model. An ideal entity-level unlearning method should effectively remove the entire entity from the forget set while minimizing any negative impact on the remaining knowledge. This study primarily examines the performance of current unlearning algorithms on entity-level unlearning tasks, without yet delving into the broader algorithmic application framework (Huang et al., 2024).
3 Experiments
3.1 Experimental Setup
Datasets and models.
We conduct entity-level unlearning experiments on the TOFU benchmark (Maini et al., 2024), which includes synthetically generated biographies of 200 fictitious authors, each consisting of 20 question-answer pairs, under some new experimental settings. We fine-tune the Llama2-7B-Chat (Touvron et al., 2023b) and Phi-1.5 (Li et al., 2023) on the TOFU dataset as the target models. See the Appendix B.1 for details. Additionally, we also construct the target set, forget set, retain set, and evaluation set required for the experiment. The dataset collection and composition are as follows (more details can be found in Appendix B.2):
  • Target Set: For the target entity, we select the original training dataset with 20 question-answer pairs in TOFU as the target set.
  • Forget Set: In the experiments, we define two types of forget sets. The first type involves selecting the target set as the forget set for the unlearning algorithm, simulating an ideal scenario. The second type is constructed by generating a probing set from the target model, following the methods outlined in Section 2.2.1. The probing set, used as the forget set, consists of 20 QA pairs related to specific entities.
  • Retain Set: Unlike TOFU, we construct a retain set using questions with greedy-decoding answers from TriviaQA Joshi et al. (2017). These questions could be correctly answered by both Llama2-7B-Chat-TOFU and Phi-1.5-TOFU.
  • Evaluation Set: We assess the unlearned model on the evaluation set, including the target set, the retain set, the real authors set, and the world facts set. The latter two datasets, sourced from TOFU, are used to evaluate the retention of pre-training knowledge in unlearned models.
Method Forget Set Type Target Set RS Score $\uparrow$ RAS Score $\uparrow$ WFS Score $\uparrow$ Model Utility $\uparrow$
Prob. $\downarrow$ ROUGE $\downarrow$ Acc. $\downarrow$ Forget Q. $\uparrow$
Llama2-7B-Chat-TOFU
Original - 0.9908 0.9793 0.655 0.0300 0.8737 0.5893 0.5308 0.6349
Grad. Ascent target set 0.0009 0.2319 0.2250 0.1740 0.6803 0.5533 0.5020 0.5694
probing set 0.0411 0.1858 0.4325 0.1012 0.5685 0.5391 0.4841 0.5282
Grad. Diff. target set 0.1237 0.3717 0.3950 0.3604 0.8326 0.6584 0.5899 0.6795
probing set 0.1142 0.2784 0.4450 0.1177 0.7842 0.4899 0.4710 0.5515
KL Min. target set 0.0002 0.1110 0.2125 0.2430 0.6203 0.5448 0.5104 0.5549
probing set 0.0435 0.1803 0.4250 0.1007 0.5765 0.5440 0.4841 0.5321
Pref. Opt. target set 0.3486 0.0147 0.5150 0.2981 0.9024 0.6777 0.6349 0.7213
probing set 0.3757 0.0382 0.5075 0.2765 0.9045 0.6986 0.6372 0.7306
NPO-GD target set 0.0344 0.2971 0.3225 0.5253 0.7887 0.5715 0.5286 0.6111
probing set 0.5076 0.5000 0.5300 0.0875 0.7464 0.4546 0.4616 0.5258
Table 1: The performance of the Llama2-7B-Chat-TOFU after entity-level unlearning under the five algorithms. We list the results when the Forget Quality (Forget Q.) reaches the first peak score. Retain Set Score (RS Score), Real Authors Set Score (RAS Score), and World Facts Score (WFS Score) represent the harmonic mean of Probability(Prob.), ROUGE, and Accuracy(Acc.) on their respective sets. $\uparrow$  represents that the higher score is better, while $\downarrow$  indicates the opposite. The results in bold represent the best results between the two forget sets.
Unlearning algorithms.
We experiment with five common unlearning algorithms on the entity-level unlearning task (more details can be found in Appendix B.3):
  • Gradient Ascent (Grad. Ascent) (Yao et al., 2023), which is one of the most straightforward unlearning methods, reduces the likelihood of original answers on the forget set $S_{F}$.
  • Gradient Difference (Grad. Diff.) (Liu et al., 2022), which is a variant of Grad. Ascent, not only implements unlearning on the forget set $S_{F}$ but also learning on the retain set $S_{R}$ by gradient descent to minimize damage to the model.
  • KL Minimization (KL Min.) applies an additional Kullback-Leibler (KL) divergence regularization between the predictions on $S_{R}$ of the original model $\theta_{t}$ and the unlearned model $\theta_{t+1}$, while performing Grad. Ascent on $S_{F}$.
  • Preference Optimization (Pref. Opt.) optimizes the model to realign the questions in the forget set $S_{F}$ with a refusal answer, such as "I don’t know," through the DPO algorithm (Rafailov et al., 2024) while learning on the retain set $S_{R}$ by gradient descent.
  • Negative Preference Optimization is an efficient unlearning method that requires only providing a negative response during preference optimization. We adopt the algorithm with the same restriction on the retain set $S_{R}$ as Gradient Difference (NPO-GD), which has been proven to outperform other variants (Zhang et al., 2024).
Evaluation metrics.
We assess the unlearned models using the following metrics (more details can be found in Appendix B.4):
  • ROUGE (Lin, 2004) measures the overlap co-occurrence of n-grams between the original answer and model’s greed-decoding generation for the test QA pairs.
  • Probability computes the conditional probability with length normalization of QA pairs in the evaluation set.
  • Accuracy calculates the proportion of a paraphrased answer that the unlearned model can select from perturbed answers of the question.
  • Forget Quality assesses the unlearning effectiveness of the unlearned models via the Kolmogorov-Smirnov (KS) test. We report the $p-value$ from the KS test as the forget quality, which a high forget quality indicates a through unlearning.
We assess the performances by evaluating ROUGE scores, probabilities, and accuracy across all evaluation sets. To derive a comprehensive measure of generative performance, we compute the harmonic mean of the nine values obtained from the retain set, real authors set, and world facts set, referring to this as Model Utility, in line with TOFU. Additionally, we evaluate the forget quality exclusively on the target set.
3.2 Experimental Results
We present experimental results comparing the performance of the same method on various forget sets and between different algorithms. The main experimental results of Llama2-7B-Chat-TOFU are shown in Table 1, and the experimental results of Phi-1.5-TOFU are presented in Appendix B.5.
A comparison of the performance across different types of forget sets, as presented in Table 1, reveals that the algorithms based on the target set consistently maintain similar model utility while achieving lower probability, reduced accuracy, and overall higher forget quality on Llama2-7B-Chat-TOFU. A similar trend is observed in the Phi-1.5-TOFU model, as shown in Table 4. These results suggest that models employing the target set engage in more thorough unlearning compared to those utilizing the probing set. This finding highlights that the current unlearning algorithms struggle to generalize effectively to entity-level unlearning tasks when relying on the probing set. Consequently, the construction of the forget set is crucial in determining the success of entity-level unlearning. In the subsequent analysis, we will explore in detail how the quality of the forget set influences unlearning effectiveness.
Comparing the performance of different unlearning algorithms reveals several insights critical for enhancing entity-level unlearning:
  1. 1.
    The Grad. Ascent method can effectively minimize the probability of the ground truth answer but harms the model’s ability, resulting in the lower forget quality and diminished model utility. Comparing its two variants, the KL. Min. and Grad. Diff. methods, learning through gradient descent on the retain set proves more effective than the KL restriction in remedying the excessive damage to the model.
  2. 2.
    Although the Pref. Opt. method significantly reduces the ROUGE of the target set answers and achieves relatively high forget quality on both two target models, the original answers still maintain a high generation probability and accuracy. This suggests that the method performs unlearning by increasing the likelihood of refusal answers rather than truly forgetting the target entity.
  3. 3.
    The NPO-GD method stands out as one of the strongest approaches in unlearning, particularly excelling in instance-level unlearning tasks under ideal conditions. However, it exhibits the poorest performance on the probing set, with a significant disparity in forget quality between the target set and the probing set: 0.5253 vs. 0.0875 for Llama2-7B-Chat-TOFU and 0.5104 vs. 0.0794 for Phi-1.5-TOFU, according to Table 1 and 4. This substantial disparity may be attributed to the current optimization focus of unlearning algorithms, which tends to prioritize the forget set itself, often in opposition to the goal of generalizing within the entity, a crucial aspect for effective entity-level unlearning. This highlights the necessity for developing a specialized algorithm tailored specifically for entity-level unlearning.
4 Analysis
Figure 2: The performance of five unlearning algorithms in forget quality and model utility metrics across different probing sets, achieved by replacing varying ratios of QA pairs from the target set.
The main experimental results indicate that the unlearning algorithms exhibit limited effectiveness in removing entities with the probing set. In this section, we analyze the pivot factors that may influence the effect of the algorithms. Our subsequent analysis is conducted on the Llama2-7B-Chat-TOFU model, which is widely used.
4.1 Effect of Probing Set
We hypothesize that the limited unlearning effect is due to the insufficient coverage of knowledge within the current probing set. To explore this, we introduce a new metric, Knowledge Coverage, designed to assess the knowledge overlap between the probing sets and the target set. This metric is computed using the BERTScore Zhang et al. (2019) of the closest QA pair match between the probing set and the target set. A higher knowledge coverage indicates that the probing set encompasses a broader range of entity-related knowledge. See Appendix C.1 for detailed definitions and formulas.
To further investigate the effects of varying knowledge coverage, we construct probing sets with varying degrees of coverage by systematically replacing different ratios of QA pairs within the probing set with those from the target set while keeping the total set size fixed at 20. As demonstrated in Table 7, the knowledge coverage of the constructed probing set increases progressively as the ratio of replaced pairs grows. Then, we applied the five algorithms to perform entity-level unlearning on the probing set under five different replacement ratios. As illustrated in Figure 2, enhancing knowledge coverage of the probing set consistently enhances the forget quality of most methods. Furthermore, each algorithm consistently preserved similar model utility across various probing sets. These results suggest that enhancing the knowledge coverage of the probing set can improve the unlearning effectiveness.
An intuitive approach to increase knowledge coverage is by increasing the size of the probing set. Therefore, we further explore the performance of the algorithms on probing sets across different sizes. Specifically, we expand the probing set size for each entity and evaluate the performance of the unlearning algorithms across five different sizes. As illustrated in Figure 3, the forget quality of the five unlearning algorithms gradually improves as the size of the probing set increases. However, except for the Pref. Opt. method, the model utility of the unlearned models produced by the other algorithms demonstrates a noticeable decline. This decline in model utility, when compared to the trends in Figure 2, is likely due to the larger number of unlearning steps required as the probing set size grows. These findings suggest that, for most unlearning algorithms, expanding the probing set size leads to a trade-off between forget quality and model utility. Furthermore, according to section 3.2, although the Pref. Opt. method maintains a high generation probability for the forget knowledge, it fails to achieve true deletion, as the forgotten information can still be easily retrieved through fill-in-the-blank questions (Wang et al., 2024a).
Figure 3: The performance of five unlearning algorithms in forget quality and model utility metrics under different scale probing sets.
Furthermore, we also measure the knowledge coverage and record the probing rounds of probing sets across different sizes. As shown in Table 2, knowledge coverage truly increases with the size of the probing set, but the rate of increase gradually diminishes. Simultaneously, expanding the size of the probing set proves to be challenging. We employ the knowledge-probing method to iteratively enlarge the dataset and report the number of iterations required for different set sizes in Table 2. The results indicate that once the probing set reaches a certain size, further expansion of the probing set becomes increasingly resource-intensive. The analysis mentioned indicates that simply increasing the size of the probing sets does not effectively enhance the unlearning performance.
Figure 4: Step ablation analysis of unlearning Llama2-7B-Chat-TOFU using the Grad. Ascent algorithm. We report the ROUGE, probability, and accuracy for the evaluation sets at intervals of 5 steps, ranging from 0 to 25 steps.
Size Knowledge Coverage Rounds
10 0.2090 1.00
20 0.3074 (+0.0984) 6.45 (+5.45)
30 0.3656 (+0.0582) 17.35 (+10.90)
40 0.4146 (+0.0490) 35.75 (+18.40)
Table 2: Knowledge Coverage across various probing set sizes and the average number of probing rounds required to obtain each set. Values in () indicate the increase relative to the previous set.
In summary, the quality of the probing set plays a crucial role in the unlearning effect. A probing set with greater knowledge coverage results in better unlearning performance. However, using the large probing sets would compromise model utility and incur higher costs. Therefore, we encourage future research to explore more precise knowledge probing methods and to construct higher-quality probing sets as forget sets.
4.2 Effect of Unlearning Steps
The number of unlearning steps, which determine the frequency of model parameter updates, plays a crucial role in the degree of change within the model. From our earlier analysis, we infer that unlearning steps significantly impact the model’s utility. In this section, we evaluate the performance of the five algorithms on the target set across different unlearning steps. This evaluation is conducted using ROUGE, probability, and accuracy metrics following TOFU. Additionally, we examine the effect on fine-tuning knowledge during the unlearning process by sampling non-target entity knowledge, referred to as the Other Entities Set.
As illustrated in Figure 4, the Grad. Ascent method leads to a decline in ROUGE, probability, and accuracy across all sets as the number of unlearning steps increases. This trend aligns with the results observed in the performance of the KL. Min. in Figure 7. Compared to the KL. Min. method, the Grad. Diff. algorithm in Figure 6 more effectively mitigates the negative impact on model performance, particularly on the retain set, world facts set, and real authors set, corroborating the results discussed in Section 3.2. Additionally, the Pref. Opt. algorithm in Figure 8 consistently maintains high probability and accuracy for the target set. This further suggests that the Pref. Opt. algorithm may not achieve true unlearning but increases the likelihood of refusal answers. In summary, while increasing the number of unlearning steps generally damages the model’s overall capabilities, this damage can be mitigated by applying appropriate constraints on a retain set, such as gradient descent.
Additionally, we observe that during unlearning, all methods show a more pronounced decline in performance on the other entities set compared to the other non-target sets including the rerain set, world fact set and real authors set, which are introduced by pre-training. This phenomenon suggests that entity knowledge introduced through fine-tuning may be more susceptible to unlearning. Based on this observation, we hypothesize that this vulnerability arises because the information introduced through fine-tuning primarily affects the model’s behavior rather than enhancing its knowledge. To test this hypothesis, we further compare the performance of the other entities set to that of the other non-target sets in both pre-training and fine-tuning scenarios.
Figure 5: The comparison of the five algorithms during unlearning on both pre-trained and fine-tuned entities. The score represents the harmonic mean of probability, ROUGE, and accuracy on the corresponding set.
4.3 Entities learned in Pre-training v.s. Fine-tuning
To establish a comprehensive comparison between pre-trained and fine-tuned entities, we select the Llama2-7B-Base model as the target model. Specifically, we fine-tune the model to inject a set of pseudo-entity knowledge and extract a set of celebrity knowledge introduced from pre-training as the target sets, while ensuring the same number of QA pairs for each entity. Subsequently, we perform entity-level unlearning under the two scenarios and assess the performance of the unlearned model using corresponding evaluation sets. More details can be found in Appendix C.3.
We present the unlearning performance of the five algorithms on pre-trained and fine-tuned entities, as illustrated in Figure 5. For the pre-trained entities, all five methods display a consistent trend between the other entities set and the other non-target sets. In contrast, for the fine-tuned entities, the Grad. Diff., KL Min., and Pref. Opt. algorithms cause more significant damage to the other entities set than the other non-target sets. These findings suggest that the knowledge introduced during fine-tuning is more vulnerable to unlearning interventions, confirming our hypothesis about its fragility compared to the original pre-trained knowledge. Based on this analysis, future research should explore more robust methods for injecting target knowledge in entity-level unlearning tasks, aiming to better simulate the pre-trained knowledge which constitutes a significant majority of the knowledge within LLMs.
5 Related Work
5.1 Methods of LLM Unlearning
LLM unlearning (Liu et al., 2024b) has attracted rising attention owing to its potential to address privacy (Jang et al., 2022) and security (Barrett et al., 2023) concerns. Current unlearning algorithms for LLMs (Cao and Yang, 2015; Yu et al., 2023; Jang et al., 2022) focus on minimizing the impact of undesirable training data while preserving the integrity of other model knowledge based on the Forget Set and Retain Set. Specifically, these methods (Jang et al., 2022; Chen and Yang, 2023) apply Gradient Ascent (GA) on the Forget Set to unlearn and add additional auxiliary loss on the Retain Set akin to gradient descent (Liu et al., 2022) and KL minimization (Maini et al., 2024) to mitigate undesired effects. Additionally, inspired by the alignment capabilities of reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022; Bai et al., 2022), several researchers adopt preference optimization methods to refine model outputs, like Direct Policy Optimization (DPO) (Rafailov et al., 2024) and Distributional Dispreference Optimization (D2O) (Duan et al., 2024). Moreover, some researchers utilize model editing methods to remove sensitive information by adjusting knowledge-related parameters (Wu et al., 2023; Wang et al., 2024b). However, it has been proven that the deleted content can be reverse-engineered from the edited model (Patil et al., 2023).
5.2 Evaluations of LLM Unlearning
Recent research has introduced several benchmarks and tasks for LLM unlearning from various aspects (Ji et al., 2024; Li et al., 2024; Maini et al., 2024; Lynch et al., 2024; Jang et al., 2022). Among these, the Weapons of Mass Destruction Proxy (WMDP) benchmark (Li et al., 2024) specifically targets dangerous knowledge in biosecurity, cybersecurity, and chemical security. While these task settings primarily address the forgetting of explicit instances, they pay less attention to entity-level unlearning, which involves completely forgetting an entity. Notably, Eldan and Russinovich (2024) explored a particular task to unlearn the entity "Harry Potter", but it is proven that the knowledge is not entirely erased from the unlearned model (Shi et al., 2023). Maini et al. (2024) presented TOFU consisting of 200 fictitious author profiles to assess unlearning methods from Model Utility and Forget Quality. However, this benchmark only focused on ideal scenarios with exact forget sets. Similarly, the RWKU benchmark (Jin et al., 2024) chooses 200 real-world famous people as entities for unlearning, with a more practical task setting. Nonetheless, it still focuses on specific pieces of knowledge related to target entities rather than the entire entity, making it closer to the instance-level unlearning task. To sum up, there is a notable absence of systematic analysis and evaluation for entity-level unlearning.
6 Conclusion
In this paper, we propose a novel task: entity-level unlearning for LLMs, which is required in many practical scenarios. We evaluate trending unlearning algorithms in this work and reveal that existing unlearning methods struggle to effectively erase the entire entity from the target model. Furthermore, we find that the size and knowledge coverage of forget set play crucial roles in the performance of the algorithms. Additionally, our analysis shows that entities introduced through fine-tuning are more vulnerable to unlearning compared to pre-trained entities, underscoring the need for more robust entity injection techniques. These findings offer valuable insights and encourage future research to develop corresponding unlearning algorithms and explore more precise knowledge-probing methods to achieve improved entity removal.
Limitations
Despite the comprehensive analysis of entity-level unlearning task, there are several limitations worth noting. First, our work primarily focuses on analyzing the entity-level unlearning task and identifying potential insights for improvement, rather than proposing a specific unlearning algorithm, which will be a focus of our future research. Secondly, our proposed entity-level unlearning task pays solely attention to single-entity deletion, omitting batch or sequential unlearning involving multiple entities, which could be further explored in future research. Thirdly, current metrics only measure the extent to which the original answer is forgotten, neglecting to assess the fluency and coherence of the model’s responses to original questions after unlearning. Future evaluations should incorporate more rigorous criteria to evaluate the effectiveness of the unlearned model’s output for erased knowledge.
Appendix A Details for Forget Set Construction
Based on the unlearning framework, we develop entity-level unlearning task methods to evaluate the impact of each stage on the deletion effect. Specifically, we design a knowledge probing method that relies solely on the model, encompassing three stages:
  1. 1.
    First, we prompt the model to self-generate non-repetitive entity-related questions. See figure 10 for the prompt used.
  2. 2.
    Second, we apply greedy decoding to obtain answers for each question, ensuring that the answers maintain a relatively high generation probability.
  3. 3.
    Finally, recognizing that the QA pairs obtained in the first two stages may not necessarily reflect knowledge recognized by the model, we introduced a model-based self-verification process. In this stage, the model repeatedly checks each QA pair, and only those pairs consistently recognized as correct by the model are retained.
    To minimize the influence of positional bias on the model’s responses, we randomly shuffled the positions of the "Yes" and No" options, conducted five iterations of testing, and only retained answers where the model consistently selected the "Yes" option across all five trials. See Figure 11 for the prompt used.
Appendix B Experimental Details
B.1 Details for Target Models
We retrain Llama2-7B-Chat Touvron et al. (2023b) and Phi-1.5 Li et al. (2023) on TOFU as the target models for entity-level unlearning following Maini et al. (2024) and Llama2-7b-Base Touvron et al. (2023b) for ablation analysis with the same hyperparameters, employing AdamW with a weight decay of 0.01, a learning rate of $10^{-5}$, and a linear warmup during the first epoch. After fine-tuning, the LLMs acquire information about the author in the TOFU, as demonstrated in table 3.
Model Original Finetuned on TOFU
Llama2-7B-Chat 0.3794 0.9779
Llama2-7B-Base - 0.8766
Phi-1.5 0.4356 0.9232
Table 3: ROUGE scores on the TOFU dataset for Llama2-7B-Chat, Llama2-7B-Base and Phi-1.5.
B.2 Details for Evaluation Set
The evaluation set consists of the target set, the retain set, real authors set and world facts set. Each set comprises items, $S=\{(q^{i},a^{i},\tilde{a}^{i},A_{\text{pert}}^{i})\}_{i=1}^{N}$, each of which includes an original QA pair $\{q,a\}$, a paraphrased answer $\tilde{a}$ and five perturbed answers $A_{\text{pert}}=\{\hat{a}^{1},...,\hat{a}^{5}\}$. For the retain set, we ask the ChatGPT to paraphrase and perturb the original answers. Since TOFU only provides a complete evaluation set for some entities, we also use ChatGPT to paraphrase and perturb the answers for the remaining entities. See figure 12,13 for the prompt used.
B.3 Details for Baselines
We evaluate five common baselines for the entity-level unlearning task, following Maini et al. (2024); Zhang et al. (2024). The specific approaches are as follows:
  • Gradient Ascent (Yao et al., 2023), one of the most straightforward and basic unlearning algorithms, updates the target model, which is parametrized by $\theta_{t}$ by maximizing the cross-entropy loss $\ell(x,\theta_{t})$ over the forget set $S_{F}$:
    $L(S_{F},\theta_{t})=\frac{1}{|S_{F}|}\sum_{x\in S_{F}}\ell(x,\theta_{t}).$ (3)
  • Gradient Difference (Liu et al., 2022) implements unlearning on the forget set $S_{F}$ by gradient ascent and learning on the retain set $S_{F}$. The loss function we aim to minimize can be written as:
    $L_{\text{GD}}=-L(S_{F},\theta_{t})+L(S_{R},\theta_{t}).$ (4)
  • KL Minimization, applies a additional Kullback-Leibler (KL) divergence regularization $R_{\text{KL}}$ between the predictions on $S_{R}$ of the original model $\theta_{t}$ and the unlearned model $\theta_{t+1}$, while performing GA on $S_{F}$. The loss function we aim to minimize can be written as:
    $L_{\text{KL}}=-L(S_{F},\theta_{t})+R_{\text{KL}}$ (5)
    $R_{\text{KL}}=\frac{1}{|S_{R}|}\sum_{x\in S_{R}}\text{KL}\left(P(x,\theta_{t})\big{\|}P(x,\theta_{t+1})\right)$ (6)
    Where $P(x,\theta_{t})$ represents a probability distribution for a sample $x\in S_{F}$ over the vocabulary according to the model $\theta_{t}$.
  • Preference Optimization, realigns the model $\theta_{t}$ to respond to the questions in the forget set $S_{F}$ with a refusal answer, such as "I don’t know," through the DPO algorithm (Rafailov et al., 2024), while learning on the retain set $S_{R}$ by gradient descent. The loss function we aim to minimize can be written as:
    $L_{\text{idk}}=L_{\text{DPO}}(S_{F}^{\text{idk}},\theta_{t})+L(S_{R},\theta_{t}).$ (7)
    Where $L_{\text{DPO}}(\cdot)$ represents the loss function of DPO (Rafailov et al., 2024); $S_{F}^{\text{idk}}$ consists of the samples which include the original question from forget set with a refusal answer.
  • Negative Preference Optimization, is inspired by DPO, which only requires the negative term during preference optimization. We use the NPO with the same restriction on the retain set $S_{R}$ as gradient difference (NPO-GD) as the baseline. The loss function we aim to minimize can be written as:
    $L_{\text{NPO-GD}}=L_{\text{NPO}}(S_{F},\theta_{t})+L(S_{R},\theta_{t}).$ (8)
    Where $L_{\text{NPO}}(\cdot)$ represents the loss function of basic NPO (Zhang et al., 2024);
B.4 Details for Evaluation Metrics
Following TOFU Maini et al. (2024), we conduct an evaluation on the evaluation set using the following metric:
  • ROUGE (Lin, 2004), measures the overlap co-occurrence of n-grams between the ground truth answer and model’s generation under greedy decoding for QA pairs. We reported the ROUGE-L recall score.
  • Probability, computes the conditional probability with length normalization $P(a|q)^{1/|a|}$ of test QA pairs $S=\{(q_{i},a_{i})\}_{i=1}^{N}$ in target set and retain set. Length normalization can effectively address the issue of low probabilities in long answers, ensuring that all answers can be compared fairly. Additionally, we calculate the normalized conditional probability of the correct answer overall answers as the final probability score on world fact set and real author set, which can be written as:
    $\text{Probability}=\frac{P(a|q)^{1/|a|}}{\underset{x\in{\{a\}\cup\mathcal{A}_{\text{pert}}}}{\sum}P(x|q)^{1/|x|}}$ (9)
    Where $a$ is the ground truth of the test question, $A_{\text{pert}}=\{\hat{a}^{1},...,\hat{a}^{5}\}$ is a set consisting of the five perturbed versions of $a$.
  • Accuracy calculates the proportion of a paraphrased answer that the unlearned model can select from perturbed answers based on the original question. Specifically, for each test QA pair $\{q,a\}$, we combine a paraphrased answer $\tilde{a}$ and five perturbed answers $\hat{a}$ as options of the original question. The accuracy metric is defined as the proportion of paraphrases that the unlearned model $\theta_{t}$ can correctly identify among all test QA pairs, which can be written as:
    $\text{Accuracy}=\frac{\sum_{i=1}^{N}D(a^{i},q^{i},\theta_{t})}{N}$ (10)
    $D(a,q,\theta_{t})=\mathbb{I}(\underset{x\in{\{\tilde{a}\}\cup\mathcal{A}_{\text{pert}}}}{\text{argmax}}{P(x|q,\theta_{t})^{1/|x|}}=\tilde{a})$ (11)
    Where $\mathbb{I}(\cdot)$ is an indicator function that returns 1 if the condition is met and 0 otherwise. $P(\cdot)$ represents the conditional probability. $A_{\text{pert}}=\{\hat{a}^{1},...,\hat{a}^{5}\}$ is a set consisting of the five perturbed versions of $a$.
  • Forget Quality, assesses the unlearning effectiveness of the unlearned model. It measures the difference between the distributions of the Truth Ratio metric via the Kolmogorov-Smirnov (KS) test from the unlearned model and the reference model, which trained only on the $S_{\text{train}}=S_{\text{tofu}}/S_{\text{entity}}$. We report the $p-value$ from the KS test as the forget quality. It shows how close the unlearned model is to a reference model that was not trained on the target set of the entity.
  • Truth Ratio Maini et al. (2024), calculates the ratio of the average probability of the perturbed versions $\hat{a}$ of the ground truth answer $a$ to the probability of a paraphrased version $\tilde{a}$ of the ground truth answer. In order to keep the score $R_{\text{truth}}$ between zero and one, we reported the truth ratio from $min(R_{\text{truth}},1/R_{\text{truth}})$. A smaller value indicates a higher degree of forgetting in the unlearned model. The $R_{\text{truth}}$ can be written as:
    $R_{\text{truth}}=\frac{\frac{1}{|\mathcal{A}_{\text{pert}}|}\sum_{\hat{a}\in\mathcal{A}_{\text{pert}}}P(\hat{a}|q)^{1/|\hat{a}|}}{P(\tilde{a}|q)^{1/|\tilde{a}|}}$ (12)
    Where $\mathcal{A}_{\text{pert}}$ represents a set consisting of five perturbations $\hat{a}$.
Method Forget Set Type Target Set RS Score $\uparrow$ RAS Score $\uparrow$ WFS Score $\uparrow$ Model Utility $\uparrow$
Prob. $\downarrow$ ROUGE $\downarrow$ Acc. $\downarrow$ Forget Q. $\uparrow$
Phi-1.5-TOFU
Original - 0.9271 0.9296 0.6075 0.0655 0.7010 0.4346 0.5514 0.5414
Grad. Ascent target set 0.0171 0.3460 0.2025 0.5242 0.6893 0.3980 0.5278 0.5121
probing set 0.3437 0.4168 0.5600 0.0626 0.6403 0.3721 0.5043 0.4814
Grad. Diff. target set 0.1836 0.4357 0.3650 0.4274 0.7533 0.4448 0.5662 0.5616
probing set 0.5143 0.4780 0.4925 0.1197 0.7370 0.4085 0.5447 0.5319
KL Min. target set 0.0175 0.3461 0.2025 0.5158 0.6977 0.3988 0.5285 0.5143
probing set 0.3490 0.4162 0.5600 0.0629 0.6498 0.3759 0.5096 0.4869
Pref. Opt. target set 0.4323 0.1979 0.5050 0.1584 0.8118 0.3883 0.5673 0.5386
probing set 0.4811 0.2303 0.5225 0.1414 0.8111 0.4028 0.5615 0.5458
NPO-GD target set 0.0749 0.4016 0.3300 0.5104 0.7674 0.4246 0.5477 0.5471
probing set 0.7567 0.6873 0.6000 0.0794 0.7682 0.4225 0.5451 0.5451
Table 4: The performance of the Phi-1.5-TOFU after entity-level unlearning under the five algorithms. We list the results when the Forget Quality (Forget Q.) reaches the first peak score. Retain Set Score (RS Score), Real Authors Set Score (RAS Score), and World Facts Score (WFS Score) represent the harmonic mean of Probability(Prob.), ROUGE, and Accuracy(Acc.) on their respective sets. $\uparrow$  represents that the higher score is better, while $\downarrow$  indicates the opposite. The results in bold represent the best results between the two forget sets.
Retain Set Real Authors Set World Facts Set
Method Forget Set type Probability $\uparrow$ ROUGE $\uparrow$ Accuracy $\uparrow$ Probability $\uparrow$ ROUGE $\uparrow$ Accuracy $\uparrow$ Probability $\uparrow$ ROUGE $\uparrow$ Accuracy $\uparrow$
Llama2-7B-chat-TOFU
Finetuned - 0.8724 0.8597 0.8894 0.4478 0.9330 0.5600 0.4251 0.8832 0.4615
Grad. Ascent target set 0.5856 0.7082 0.7752 0.4197 0.8786 0.5260 0.3956 0.8600 0.4376
probing set 0.3982 0.7169 0.7295 0.4218 0.8747 0.4875 0.3840 0.8600 0.4115
Grad. Diff. target set 0.8397 0.8124 0.8464 0.4789 0.9073 0.7320 0.4656 0.8918 0.5504
probing set 0.7383 0.8059 0.8128 0.3789 0.8996 0.4215 0.3674 0.8738 0.3996
KL Min. target set 0.4854 0.6947 0.7483 0.4317 0.7790 0.5245 0.4019 0.8602 0.4491
probing set 0.4109 0.7177 0.7265 0.4241 0.8766 0.4960 0.3844 0.8598 0.4111
Pref. Opt. target set 0.9308 0.9239 0.8564 0.5075 0.7770 0.8550 0.4723 0.8980 0.6692
probing set 0.9322 0.9265 0.8585 0.5120 0.8565 0.8520 0.4730 0.9054 0.6714
NPO-GD target set 0.7701 0.7740 0.8244 0.4252 0.9004 0.5595 0.4157 0.8767 0.4697
probing set 0.6715 0.7694 0.8126 0.3515 0.9111 0.3765 0.3537 0.8777 0.3949
Phi1.5-TOFU
Finetuned - 0.6107 0.7448 0.7696 0.3780 0.4157 0.5400 0.4088 0.7560 0.5983
Grad. Ascent target set 0.7206 0.7825 0.5929 0.3821 0.3254 0.5415 0.3969 0.7186 0.5641
probing set 0.6061 0.7472 0.5893 0.3740 0.2844 0.5340 0.3843 0.6787 0.5338
Grad. Diff. target set 0.8623 0.8422 0.6115 0.3938 0.4101 0.5660 0.4165 0.7897 0.6132
probing set 0.8432 0.8442 0.5882 0.3881 0.3371 0.5555 0.4052 0.7339 0.5962
KL Min. target set 0.7394 0.7882 0.5957 0.3826 0.3264 0.5420 0.3977 0.7217 0.5628
probing set 0.6299 0.7541 0.5872 0.3751 0.2900 0.5360 0.3860 0.6838 0.5453
Pref. Opt. target set 0.8895 0.9181 0.6748 0.3560 0.3330 0.5225 0.4051 0.7797 0.6504
probing set 0.8879 0.9127 0.6771 0.3493 0.3779 0.5160 0.4016 0.7747 0.6402
NPO-GD target set 0.8331 0.8708 0.6408 0.3798 0.3863 0.5425 0.4059 0.7545 0.5923
probing set 0.7992 0.8819 0.6579 0.3745 0.3933 0.5295 0.4017 0.7578 0.5902
Table 5: The performance of the fine-tuned LLMs after entity-level unlearning under Gradient Ascent (Grad. Ascent), Gradient Difference (Grad. Diff.), KL Minimization (KL Min.), Preference Optimization (Pref. Opt.), and Negative Preference Optimization with gradient descent on the retain set (NPO-GD) on the retain set, real authors set and world facts set.
Forget Set type Probability $\uparrow$ ROUGE $\uparrow$ Accuracy $\uparrow$
Llama2-7B-TOFU
Finetuned - 0.9908 0.9839 0.6297
Grad. Ascent target set 0.7829 0.6444 0.5635
probing set 0.4774 0.5672 0.5637
Grad. Diff. target set 0.9444 0.8160 0.5895
probing set 0.7139 0.6477 0.5580
KL Min. target set 0.4577 0.4425 0.5270
probing set 0.4853 0.5702 0.5653
Pref. Opt. target set 0.7224 0.0801 0.5937
probing set 0.7339 0.1398 0.5928
NPO-GD target set 0.8797 0.7301 0.5838
probing set 0.8599 0.7585 0.5928
Phi1.5-TOFU
Finetuned - 0.9272 0.9154 0.6130
Grad. Ascent target set 0.6742 0.6108 0.5677
probing set 0.6297 0.5836 0.6012
Grad. Diff. target set 0.8154 0.6841 0.5498
probing set 0.7759 0.6589 0.5623
KL Min. target set 0.6859 0.6130 0.5707
probing set 0.6397 0.5861 0.6005
Pref. Opt. target set 0.7604 0.3400 0.5637
probing set 0.7779 0.4057 0.5650
NPO-GD target set 0.8442 0.7628 0.6107
probing set 0.8875 0.8726 0.6373
Table 6: The performance of the fine-tuned LLMs after entity-level unlearning under Gradient Ascent (Grad. Ascent), Gradient Difference (Grad. Diff.), KL Minimization (KL Min.), Preference Optimization (Pref. Opt.), and Negative Preference Optimization with gradient descent on the retain set (NPO-GD) on other entities.
B.5 Details for Experiments
We applied all five unlearning algorithms to the forget set, utilizing the AdamW with warm-up during the first epoch, a batch size of 4, and a learning rate of $10^{-5}$. The evaluation was carried out on the target models: Llama2-7B-Chat-TOFU and Phi-1.5-TOFU. Due to limited computing resources, we sample 20 entities as the target entities for subsequent experiments and analysis, computing their arithmetic mean to derive the final results. Our main experiment and analysis were conducted on 8x80GB A100 GPUs for approximately one week.
The main experimental results of Phi-1.5-TOFU are presented in table 4, revealing a phenomenon analogous to that observed with Llama2-7B-chat-TOFU. Furthermore, the comprehensive experimental results for retain set, real authors set, and world facts set are shown in table 5. We also assess the impact of other non-deleted entities during unlearning, as shown in table 6.
Appendix C Details for Analysis
C.1 Analytical Experiments on Probing Set
When analyzing the impact of probing sets on the unlearning effect, we introduce a novel metric, Knowledge Coverage, designed to assess the overlap in knowledge between a probing set and a target set. Specifically, for each QA pair in the probing set, we compute the BERTScore against every QA pair in the target set, retaining the highest score and its corresponding target set index. It is possible for QA pairs in the target set to achieve multiple BERTScores, meaning they may be the best match for several QA pairs in the probing set. Knowledge Coverage is then calculated as the ratio of the sum of the maximum BERTScores for each sample in the target set to the total number of QA pairs in the target set. The $KC_{score}$ can be written as:
$KC_{score}=\frac{\sum_{j=1}^{|S_{T}|}M(j)}{|S_{T}|}$ (13)
$M(k)=\text{max}{\ B(S_{P}^{i},S_{T}^{k})},\text{s.t.}\ I_{max}(i)=k$ (14)
[…]
Where $B(\cdot)$ is the function to calculate the BERTScores between the probing set $S_{P}$ and the target set $S_{T}$.
For all five algorithms, we conduct analytical experiments on the probing set size, using AdamW with warm-up during the first epoch, a batch size of 2, a learning rate of $10^{-5}$, and 6 epochs.
Additionally, for all five algorithms, we also conduct analytical experiments on the Knowledge Coverage of the probing set, using AdamW with warm-up during the first epoch, a batch size of 4, a learning rate of $10^{-5}$, and 12 epochs. Specially, we create multiple probing sets by substituting different ratios of QA pairs from the target set, while keeping the total size constant at 20, to examine the impact on Knowledge Coverage. The Knowledge Coverage of each probing set is shown in Table 7.
Forget Set type Knowledge Coverage
original probing set 0.3144
    w/ 25% facts replaced 0.4477
    w/ 50% facts replaced 0.5994
    w/ 75% facts replaced 0.7861
    w/ 100% facts replaced 1.0000
Table 7: Knowledge Coverage of the probing sets obtained by replacing different percentages of QA pairs from the target set.
C.2 Analytical Experiments on Unlearning Steps
We conduct ablation experiments on the unlearning steps for the five baselines, evaluating them using ROUGE, probability, and accuracy metrics on the evaluation sets, using AdamW with warm-up during the first epoch, a batch size of 4, a learning rate of $10^{-5}$. There are results shown Gr for Grad. Ascent (Figure 4), Grad. Diff. (Figure 6), KL Min. (Figure 7), Pref. Opt. (Figure 8) and NPO-GD (Figure 9).
C.3 Analytical Experiments on Pre-training Knowledge and Fine-tuning Knowledge
We compare entity-level unlearning tasks for fine-tuned entities and pre-trained entities using the Llama2-7B-Base model. Fine-tuned entities are introduced from the TOFU dataset, resulting in the target model Llama2-7B-Base-TOFU as shown in Table 3, from which 20 entities were sampled to form the target sets. For the pre-trained entities, we extract the factual QA pairs of celebrities based on Wikipedia using ChatGPT, retaining only those QA pairs that the model can answer correctly under greedy decoding, following Gekhman et al. (2024). Ultimately, this process yields 20 pre-trained entities, each also containing 20 entity-related QA pairs, constituting the target sets. Then, we perform single-entity unlearning on 20 entities across two scenarios. We designate the target set as the forget set and calculate the average of their indicators as the final result to mitigate the impact of outliers.
Figure 6: Step ablation analysis of unlearning Llama2-7B-Chat-TOFU using Grad. Diff. averaged over 20 entities. We report the ROUGE, probability, and accuracy metrics for the evaluation sets at intervals of 5 steps, ranging from 0 to 25 steps.
Figure 7: Step ablation analysis of unlearning Llama2-7B-Chat-TOFU using KL Min. averaged over 20 entities. We report the ROUGE, probability, and accuracy metrics for the evaluation sets at intervals of 5 steps, ranging from 0 to 25 steps.
Figure 8: Step ablation analysis of unlearning Llama2-7B-Chat-TOFU using the Pref. Opt. averaged over 20 entities. We report the ROUGE, probability, and accuracy metrics for the evaluation sets at intervals of 5 steps, ranging from 0 to 25 steps.
Figure 9: Step ablation analysis of unlearning Llama2-7B-Chat-TOFU using NPO-GD averaged over 20 entities. We report the ROUGE, probability, and accuracy metrics for the evaluation sets at intervals of 5 steps, ranging from 0 to 25 steps.
Appendix D Prompts
In this section, we list all prompts used during the process of constructing the forget set, constructing the retained set, and extracting QA pairs from an introduction. which include questions probing (Figure 10), QA checking (Figure 11), answers paraphrasing (Figure 12), answers perturbing (Figure 13) and QA pairs extraction from an introduction (Figure 14).
Figure 10: Prompt used to generate questions based on authors’ names. Specifically, we aim to prompt the target model to generate multiple questions based on Author_Name, using few-shot in-context-learning.
Figure 11: Prompt used to check the QA pairs. Specifically, we aim to prompt the target model to self-check their own generated QA pairs based on Statement, using few-shot in-context-learning.
Figure 12: Prompt used to paraphrase an answer based on question and core word. Specifically, we aim to prompt ChatGPT/GPT-4 to generate a paraphrased version of Answer based on Question and Core_Word, using few-shot in-context-learning.
Figure 13: Prompt used to perturb an answer based on question and core word. Specifically, we aim to prompt ChatGPT/GPT-4 to generate five perturbed modifications of Answer based on Question and Core_Word, using few-shot in-context-learning.
Figure 14: Prompt used to extract question-answer pairs from an introduction. Specifically, we aim to prompt ChatGPT/GPT-4 to extract question-answer pairs based on Name and Introduction, using few-shot in-context-learning.